OpenStudy (anonymous):

Write the equation of the circle centered at (9 , -10 ) that passes through (19,15).

OpenStudy (anonymous):

first find the distance between (9 , -10 ) and (19,15). this will give the radius

OpenStudy (anonymous):

equation of the circle centred at (h,k) and radius a unit is given by (x-h)^2 +(y-k)^2=a^2

OpenStudy (anonymous):

thats the part im having issues with... i keep getting 5\[\sqrt{13}\]

OpenStudy (anonymous):

for the distance

OpenStudy (anonymous):

have u checked ur calculation

OpenStudy (anonymous):

ya its not right... but idk what i did wrong

OpenStudy (anonymous):

i can show you what i did

OpenStudy (anonymous):

distance between (9,-10) and (19,15) =sqrt ( (19-9)^2 + (15+10)^2 )

OpenStudy (anonymous):

ok show

OpenStudy (anonymous):

\[d=\sqrt{(19-9)^2+(15+10)^2}\]

OpenStudy (anonymous):

\[\sqrt{10^2+25^2}\]

OpenStudy (anonymous):

\[\sqrt{100+225}\]

OpenStudy (anonymous):

\[\sqrt{325}\]

OpenStudy (anonymous):

correct =sqrt(100+625) so u get sqrt(725) =(sqrt(5*5*29) = 5*29

OpenStudy (anonymous):

25^2=625 but u are writing 325 as 15^2=325

OpenStudy (anonymous):

oh wow i messed up 25^2 -_-

OpenStudy (anonymous):

thx ur awesome

OpenStudy (anonymous):

so its \[5\sqrt{29}\] ?

OpenStudy (anonymous):

yup any yw

OpenStudy (anonymous):

1 more question... how do i make it into the radius?